OpenStudy (anonymous):

6x - 6y=42 7x+6y=-48

OpenStudy (anonymous):

now add both equations...

OpenStudy (anonymous):

you left with: 13x=-6 can you solve for x?

OpenStudy (anonymous):

wait does -6 equal x?

OpenStudy (anonymous):

no if you add 1 & 2 : 6x+7x = 42-48 (6y will cancel out)

OpenStudy (anonymous):

if 13x=-6 x=-6/13 right?

OpenStudy (anonymous):

sooo?.... is that a fraction -6/13 what is y?

OpenStudy (anonymous):

I NEED HELP =(

OpenStudy (anonymous):

now, you put the value of x in any original equation, let's say 1st...

OpenStudy (anonymous):

6x-6y=42 divide both sides by 6: x-y=7

OpenStudy (anonymous):

so i divide 6 by 6?

OpenStudy (anonymous):

x=-6/13,so y=x-7=-6/13 -7=-97/13

OpenStudy (anonymous):

x=6?

OpenStudy (anonymous):

this is your solution: x=-6/13 & y=-97/13

OpenStudy (anonymous):

are u sure sir?

OpenStudy (anonymous):

pretty sure...

OpenStudy (anonymous):

that was wrong i mean i got it wrong..... its supposed to be a whole number not a fraction =(

OpenStudy (anonymous):

unless there is a mistake in original problem... I tried different ways

OpenStudy (anonymous):

check the original problem... did you put it right?

OpenStudy (anonymous):

yea i did it basically like how u did the first one

OpenStudy (anonymous):

let me try again... give a min

OpenStudy (anonymous):

its to late...

OpenStudy (anonymous):

I'm getting the same numbers... sorry!

OpenStudy (anonymous):

can u help me with another please?